OpenStudy (ashy98):

An inequality is shown below: -np - 5 ≤ 4(c - 2) Which of the following solves for n?

OpenStudy (ashy98):

\[A.) n \ge-\frac{ 4c-3 }{ p }\] \[B.) n \ge-\frac{ 4c-13 }{ p }\] \[C.) n-\le-\frac{ 4c-3 }{ p }\] \[D.) n \le-\frac{ 4c-13 }{ p }\]

OpenStudy (ashy98):

@esam2 @yiffboy @wolf1728 @taylor1216